OpenStudy (anonymous):

I Need The Have The Work For 4y + 15 = 6y + 11 Asap (:

OpenStudy (saifoo.khan):

y = 2

OpenStudy (diyadiya):

6y-4y=15-11 2y=4 y=2

OpenStudy (anonymous):

Thank(: